Country A produces about five times the amount of diamonds in carats produced in country B. If the total amount produced in both countries is 24,000,000
carats, find the total produced in each country. Please help me solve this problem.
A 17 foot piece of string that is cut into two pieces so that the longer piece is five feet longer than twice the shorter piece. If the shorter piece is x feet long find the lengths of both sides

Answer by macston(5194)

.
B=amount produced Country B
A=amount produced Country A=5B
A+B=24,000,000
5B+B=24,000,000
6B=24,000,000
B=4,000,000 ANSWER 1: Country B produces 4,000,000 carats.
A=5B=5(4,000,000)=20,000,000 ANSWER 2: Country A produces 20,000,000 carats.
.
A 17 foot piece of string that is cut into two pieces so that the longer piece is five feet longer than twice the shorter piece. If the shorter piece is x feet long find the lengths of both sides
.
S=length of shorter piece
L=length of longer piece=2S+5ft
S+L=17 feet
S+(2S+5ft)=17ft
3S+5ft=17ft
3S=12ft
S=4ft ANSWER 1: The shorter piece was 4 feet.
L=2S+5ft=2(4ft)+5ft=13ft ANSWER 2: The longer piece was 13 feet.